# Subdiffusive concentration for the chemical distance in Bernoulli percolation

### v1

@@ 2.1. Effective radius and its applications
Since ${\mathrm{T}}_{x,y}$ is increasing in $t_{e}$, the first inequality in (2.5) is trivial. On the other hand, if $e\notin\gamma$ then
@@
Suppose that $e\in\gamma$ and ${\mathcal{U}}_{e}^{c}$ occurs (and thus neither $x$ nor $y$ is in $\Lambda_{3R_{e}}(e)$). Applying Proposition 2.5 to $\gamma\in\mathbb{G}_{*}(x,y)$ and $e\in\gamma$, there exists a path $\eta_{e}$ between $x$ and $y$ such that
${\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})\leq{\mathrm{T}}(\eta_{e}\setminus\gamma_{z})=|\eta_{e}\setminus\gamma_{z}|\leq C_{*}R_{e}.$
@@
We first review some intrinsic properties of chemical distance and crossing cluster of supercritical Bernoulli percolation.
@@
[Pis96, Theorem 2] There exists a constant $c>0$ such that for all $x\in{\mathbb{Z}}^{d}$ and $t>0$
$\mathbb{P}(\|x-x^{*}\|_{\infty}\geq t)\leq c^{-1}\exp(-ct^{d-1}).$ (2.6)
@@
[Gri89, Theorem 7.68] There exists a constant $c>0$ such that for all $t>0$,
@@
[AP96, (4.49)] There exist $\rho,c>0$ such that for every $x\in{\mathbb{Z}}^{d}$ and $t\geq\rho\|x\|_{\infty}$,
$\max\{\mathbb{P}(\infty>{\mathrm{D}}(0,x)\geq t),\mathbb{P}({\mathrm{D}}^{*}(0,x)\geq t)\}\leq c^{-1}\exp(-ct).$ (2.7)
@@
[Gri89, Lemma 7.104] For any $\varepsilon>0$, there exists $c>0$ such that for all $t>0$,
(two disjoint open clusters with diameter at least εt in $\mathbb{P}(\,\exists\,\text{two disjoint open clusters with diameter at least $\varepsilon t$ in $\Lambda_{t}$})\leq c^{-1}\exp(-ct).$
@@ 2.2.1. Good box
We present in this part a condition called good box under which the bypass in the definition of effective radius can be easily constructed. For $m\in{\mathbb{N}}$ and $X\subset{\mathbb{Z}}^{d}$, let ${\mathcal{B}}_{m}(X)$ denote the set of all boxes of side-length $m$ in $X$. Let $\rho$ be the constant defined as in Lemma 2.10, and set
@@
Among the properties of a good box, (iv) is highly non-trivial and is treated in the following result.
@@
For any $p>p_{c}(d)$, there exists a positive constant $c_{*}\in(0,1)$ such that for all $t\geq 1/c_{*}$ and $1\leq N\leq\exp(c_{*}\log^{2}n)$,
$\mathbb{P}({\mathcal{E}}_{N})\leq c_{*}^{-1}\exp(-c_{*}N),$
@@
${\mathcal{L}}_{N}:=\{\forall\text{ crossing cluster }{\mathcal{C}}\subset\Lambda_{2N},\,\forall\text{ open cluster }{\mathcal{D}}\subset\Lambda_{2N}\text{ with }\operatorname{Diam}({\mathcal{D}})\geq N:{\mathcal{C}}\cap{\mathcal{D}}\neq\varnothing\}.$
@@
$({\mathcal{E}}^{\prime}_{N})^{c}\cap{\mathcal{L}}_{N}\subset({\mathcal{E}}_{N})^{c},$ (2.8)
@@
By Lemma 2.11, there exists $c_{1}>0$ such that ${\mathbb{P}}({\mathcal{L}}_{N}^{c})\leq c_{1}^{-1}\exp(-c_{1}n)$, and thus (2.8) gives
$\mathbb{P}({\mathcal{E}}_{N})\leq\mathbb{P}({\mathcal{E}}^{\prime}_{N})+c_{1}^{-1}\exp(-c_{1}N).$ (2.9)
Suppose that ${\mathcal{E}}^{\prime}_{N}$ occurs, and let $\pi\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(\Lambda_{tN})$ be a path that realizes ${\mathcal{E}}^{\prime}_{N}$. We assume further that the diameter of $\pi$ is achieved in the first coordinate direction, i.e., there exist $u,v\in\pi$ with $u_{1}-v_{1}=\operatorname{Diam}(\pi)=N/2$. Let $\pi^{*}$ be the subpath of $\pi$ from $u$ to $v$. Since $\pi\in\mathbb{G}_{*}(\Lambda_{tN})$, so is $\pi^{*}$.
We claim that for any $a>0$, there exists $\varepsilon\in(0,1/3)$ such that for all $N\leq\exp(\varepsilon\log^{2}n)$ and $t\geq 1/\varepsilon$,
${\mathbb{P}}(({\mathcal{L}}_{N}^{\prime})^{c})\leq\varepsilon^{-1}\exp(-\varepsilon N),\textrm{ where }{\mathcal{L}}_{N}^{\prime}:=\{\forall\,\gamma\in\cup_{x,y\in\Lambda_{N}}\mathbb{G}(x,y;\Lambda_{tN}):|{\mathrm{clo}}(\gamma)|\leq\tfrac{aN}{\log N}\},$ (2.10)
where recall that ${\mathrm{clo}}(\gamma)=\{e\in\gamma:e\textrm{ is closed}\}$. Indeed, by Lemmas 2.8 and 2.10 with the remark that $d\geq 3$, there exists $\alpha\in(0,1)$ such that
@@
for all $1/\varepsilon\leq N\leq\exp(\varepsilon\log^{2}n)$ with $\varepsilon=\varepsilon(a,\alpha)$ sufficiently small. Hence, (2.10) follows.
@@
where recall that $u$ is the starting point of $\pi^{*}$. Notice that for all $q\in[K]$, $\pi^{*}$ has to cross the slab $S_{u,q}$ from $H_{u,q-1}$ to $H_{u,q}$. For each $q\in[K]$, let $\pi^{*}_{q}\subset\pi^{*}$ denote the connected subpath of $\pi^{*}\cap S_{u,q}$ such that $H_{u,q-1}\xleftrightarrow{\Lambda_{N}}H_{u,q}$ via $\pi^{*}_{q}$. Hence, on the event ${\mathcal{E}}^{\prime}_{N}$,
[…]
Let $I:=\{q\in[K]:\pi_{q}^{*}\in{\mathbb{O}}(\Lambda_{N})\}$ be the set of open subpaths of $\pi^{*}$. As $\pi^{*}\in\mathbb{G}_{*}(u,v;\Lambda_{tN})$, there exists $\eta\in\mathbb{G}(u,v;\Lambda_{tN})$ such that $\pi^{*}\setminus\eta$ is open. Then for any $a>0$, conditioned on ${\mathcal{L}}_{N}^{\prime}$ (defined in (2.10) with $t\geq 1/\varepsilon$ and $\varepsilon=\varepsilon(a)$ sufficiently small), we get that
$\textrm{clo}(\pi^{*})\leq\textrm{clo}(\eta)\leq\frac{aN}{\log N},$
@@
$|I|\geq K-|{\mathrm{clo}}(\pi^{*})|\geq K-\frac{aN}{\log N}\geq\frac{N}{4L},$ (2.11)
@@
[…]
where in the last line we have used the independence of events in disjoint slabs. By Lemmas 2.9, 2.11 and union bound, there exists a positive constant $C$ such that if $L\geq C\log N$,
[…]
@@
$\mathbb{P}({\mathcal{E}}^{\prime}_{n}\cap{\mathcal{L}}^{\prime}_{N})\leq\exp(-N/(8C)).$ (2.12)
Finally, we clarify the choice of parameters. We take $a=1/(10C)$ and $L=\lceil C\log N\rceil$ under which $L\leq\log N/(5a)$ and $L\geq C\log N$. Using (2.12) and (2.10),
@@
with $\varepsilon=\varepsilon(a)$ given in (2.10). Combining this with (2.9) and taking $c_{*}=\frac{1}{4}\min\{\varepsilon,c_{1},1/8C\}$, we obtain the desired result. ∎
@@
Fix $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. It follows from Lemmas 2.9 and 2.11 that
$\mathbb{P}(\Lambda_{3N}(e)\text{ does not satisfy (i)})\leq c_{1}^{-1}\exp(-c_{1}N_{\rho}),$ (2.13)
@@
[…]
$\leq\mathbb{P}(\exists\,x,y\in\Lambda_{3N}(e),\|x-y\|_{\infty}<2N_{\rho},{\mathrm{D}}(x,y)<\infty,{\mathrm{D}}(x,y)\geq 4\rho N_{\rho})\leq c_{2}^{-1}\exp(-c_{2}N).$ (2.14)
@@
$\leq c_{3}^{-1}|{\mathrm{A}_{N}(e)}|^{2}\exp(-c_{3}N_{\rho})\leq c_{3}^{-1}\exp(-c_{3}^{2}N_{\rho}).$ (2.15)
[…]
Notice that given a path $\pi\subset\Lambda_{3N}(e)$ with $\operatorname{Diam}(\pi)\geq N_{\rho}$, there exist $x,y\in\pi$ such that $\|x-y\|_{\infty}\geq N_{\rho}$. That is, there exists a sub-path $\pi^{\prime}\subset\pi_{x,y}\cap\Lambda_{N_{\rho}/4}(x)$ such that $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$, where $\pi_{x,y}$ is the sub-path of $\pi$ from $x$ to $y$. Notice that we always can find $\Lambda\in\mathcal{B}_{N_{\rho}/2}(\Lambda_{3N}(e))$ such that $\Lambda$ contains $\Lambda_{N_{\rho}/4}(x)\cap\Lambda_{3N}(e)$, and hence $\pi^{\prime}\subset\Lambda$ with $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$. Therefore, it follows from (iv’) that there exists a vertex $z\in\Lambda_{3N}(e)$, a crossing cluster ${\mathcal{C}}^{\prime}\subset\Lambda_{N_{\rho}/2}(z)\subset\Lambda_{3N}(e)$, and a sub-path $\pi^{\prime}\in{\mathcal{P}}(\Lambda_{N_{\rho}/4}(z))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N_{\rho}/4}(z))$ with $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$, such that $\pi^{\prime}\cap{\mathcal{C}}^{\prime}=\varnothing$. Using Lemma 2.14, there exists $c_{4}>0$ such that for all $N\geq 1$
[…]
@@
Combining this estimate with (2.13), (2.2.1), and (2.2.1) yields the desired result. ∎
@@
$\{\Lambda_{3N}(e)\text{ is }\textbf{ good}\}\subset\mathcal{V}^{1}_{N}(e)\cap\mathcal{V}^{2}_{N}(e).$ (2.16)
@@
Next, we assert that $\mathcal{V}^{2}_{N}(e)$ holds. For $x,y\in\Lambda_{3N}(e)$ with ${\mathrm{D}}^{\Lambda_{3N}(e)}(x,y)<\infty$, there exists an open path $\gamma\subset\Lambda_{3N}(e)$ joining $x$ to $y$. Thus, we can find a sequence $(z_{i})_{i=0}^{h}\subset\gamma$ with $h\leq(6N/N_{\rho})\leq(100\rho^{2})^{d}$ such that $z_{0}=x,z_{h}=y$ and
$\forall i\in[h],\quad\|z_{i-1}-z_{i}\|_{\infty}\leq 2N_{\rho}.$ (2.17)
Then, by Definition 2.12 (ii), ${\mathrm{D}}^{\Lambda_{4N}(e)}(y_{i-1},y_{i})\leq 4\rho N_{\rho}$. Therefore,
${\mathrm{D}}^{\Lambda_{4N}(e)}(x,y)\leq\sum_{i=1}^{h}{\mathrm{D}}^{\Lambda_{4N}(e)}(y_{i-1},y_{i})\leq h(4\rho N_{\rho})\leq C_{*}N,$
and thus $\mathcal{V}^{2}_{N}(e)$ holds. In summary, we have proved (2.16) and completed the proof of this lemma. ∎

### v2

@@ 2.1. Effective radius and its applications
Since ${\mathrm{T}}_{x,y}$ is increasing in $t_{e}$, the first inequality in (21) is obvious. Next, we consider the second inequality. Assume that $e$ is open (i.e. $t_{e}=1$). If $e\notin\gamma$, then closing $e$ ($t_{e}=\log^{2}n$) has no effect on the geodesic, and hence
@@
If $e\in\gamma$ and ${\mathcal{U}}_{e}^{c}$ occurs, then neither $x$ nor $y$ belongs to $\Lambda_{3R_{e}}(e)$. Applying Proposition 2.7 to $\gamma\in\mathbb{G}_{*}(x,y)$ and $e\in\gamma$, there exists another path $\eta_{e}$ between $x$ and $y$ such that
${\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})\leq{\mathrm{T}}(\eta_{e}\setminus\gamma)=|\eta_{e}\setminus\gamma|\leq C_{*}R_{e}.$
@@
We first review some intrinsic properties of chemical distance and crossing cluster of supercritical Bernoulli percolation. We recall that an open cluster $\mathcal{C}$} is a crossing cluster of a box $\Lambda$ if for all $d$ directions, there is an open path in ${\mathcal{C}}\cap\Lambda$ connecting the two opposite faces of $\Lambda$.
@@
[29, Theorem 2] There exists a constant $c>0$ such that for all $x\in{\mathbb{Z}}^{d}$ and $t>0$
$\mathbb{P}(\|x-x^{*}\|_{\infty}\geq t)\leq c^{-1}\exp(-ct^{d-1}).$ (22)
@@
[27, Theorem 7.68] There exists a constant $c>0$ such that for all $t>0$,
@@
[25, Corollary 2.2, Lemma 2.3] There exist constants $\rho\geq 1$ and $c>0$ such that for every $x\in{\mathbb{Z}}^{d}$ and $t\geq\rho\|x\|_{\infty}$,
$\max\{\mathbb{P}(\infty>{\mathrm{D}}(0,x)\geq t),\mathbb{P}({\mathrm{D}}^{*}(0,x)\geq t)\}\leq c^{-1}\exp(-ct).$ (23)
@@
[27, Lemma 7.104] For any $\varepsilon>0$, there exists $c>0$ such that for all $t>0$,
[…]
Given $t>\ell>0$, we define slab
$S(\ell,t):=\{x=(x_{1},\ldots,x_{d})\in\Lambda_{t}:0\leq x_{1}\leq\ell\}.$
Lemma 2.14.
There exists a constant $C>0$ such that for all $C\log t\leq\ell<t$,
$\mathbb{P}(\exists\,\gamma\subset S(\ell,t):\operatorname{Diam}(\gamma)\geq\ell,\gamma\not\xleftrightarrow{S(\ell,t)}\partial\Lambda_{t})\leq C\exp(-\ell/C).$ (24)
Proof.
We say an open cluster $\mathcal{C}$} crosses the slab $S(\ell,t)$ if for all $d$ directions, there is an open path of ${\mathcal{C}}\cap S(\ell,t)$ connecting two opposite faces of $S(\ell,t)$. Observe that (d1)(d-1) pairs of opposite faces of $S(\ell,t)$ are contained in $\partial\Lambda_{t}$ (except two opposite faces $\{x\in S(\ell,t):x_{1}=0\}$ and $\{x\in S(\ell,t):x_{1}=\ell\}$). Thus, if $\mathcal{C}$} is a crossing cluster in $S(\ell,t)$, then ${\mathcal{C}}\cap\partial\Lambda_{t}\neq\varnothing$. Let ${\mathcal{B}}_{m}(X)$ be the set of all boxes of side-length $m$ in $X\subset{\mathbb{Z}}^{d}$. Define
[…]
It follows from Lemmas 2.11 and 2.13 that there exists a constant $C_{1}>0$ such that for all $\ell\geq C_{1}\log t$,
$\mathbb{P}({\mathcal{W}}^{1}_{\ell,t})\geq 1-C_{1}\exp(-\ell/C_{1}).$
Thanks to union bound and Lemma 2.13 again, there exists a constant $C_{2}>0$ such that for all $\ell\geq C_{2}\log t$,
$\mathbb{P}({\mathcal{W}}^{2}_{\ell,t})\geq 1-C_{2}\exp(-\ell/C_{2}),$
where
[…]
Suppose that ${\mathcal{W}}^{1}_{\ell,t}\cap{\mathcal{W}}^{2}_{\ell,t}$ occurs. Then, there exists a crossing cluster $\mathcal{C}$} in $S(\ell,t)$ such that for all $\Lambda\in{\mathcal{B}}_{\ell}(S(\ell,t)$ and any open path $\eta\subset\Lambda$ with diameter at least $\ell/2$, ${\mathcal{C}}\cap\eta\neq\varnothing$. Notice that if $\gamma\subset S(\ell,t)$ with $\operatorname{Diam}(\gamma)\geq\ell$, there exist $x,y\in\gamma$ such that $\|x-y\|_{\infty}\geq\ell$. Thus, there exists a subpath $\gamma^{\prime}\subset\gamma_{x,y}\cap\Lambda_{\ell/2}(x)$ such that $\operatorname{Diam}(\gamma^{\prime})\geq\ell/2$. Moreover, there exists a box $\Lambda\in\mathcal{B}_{\ell}(S(\ell,t))$ such that $\Lambda$ contains $\Lambda_{\ell/2}(x)\cap S(\ell,t)$. Therefore, $\gamma^{\prime}\cap{\mathcal{C}}\neq\varnothing$ due to the assumption of ${\mathcal{W}}^{1}_{\ell,t}\cap{\mathcal{W}}^{2}_{\ell,t}$. Thus $\gamma\xleftrightarrow{S(\ell,t)}\partial\Lambda_{t}$ since ${\mathcal{C}}\cap\partial\Lambda_{t}\neq\varnothing$. In conclusion,
$\mathbb{P}(\forall\gamma\in S(\ell,t):\operatorname{Diam}(\gamma)\geq\ell,\gamma\xleftrightarrow{S(\ell,t)}\partial\Lambda_{t})\geq\mathbb{P}({\mathcal{W}}^{1}_{\ell,t}\cap{\mathcal{W}}^{2}_{\ell,t})\geq 1-C\exp(-\ell/C),$
with $C:=2\max\{C_{1},C_{2}\}>0$. ∎
@@ 2.2.1. Good box
We present in this part a condition called good box under which the bypass in the definition of effective radius can be easily constructed. Let $\rho$ be the constant defined as in Lemma 2.12, and set
@@
Among the properties of a good box, (iv) is the most complicated to prove and is not standard, which is treated in the following result.
@@
For any $p>p_{c}(d)$, there exists a constant $c_{*}\in(0,1)$ such that for all $t\geq 1/c_{*}$ and $1\leq N\leq\exp(c_{*}\log^{2}n)$,
$\mathbb{P}({\mathcal{E}}_{N})\leq c_{*}^{-1}\exp(-c_{*}N),$ (25)
@@
In the event ${\mathcal{E}}_{N}$, if $\pi\in{\mathcal{P}}(\Lambda_{N})\cap{\mathbb{O}}(\Lambda_{N})$, i.e., $\pi$ was opened, then (25) obviously follows from Lemma 2.13. To prove this lemma, the key point here is to show that $\pi$ does not contain many closed edges, and thus has a large number of open segments. The remainder of the proof follows by using the slab technique.
@@
[…]
@@
$({\mathcal{E}}^{\prime}_{N})^{c}\cap{\mathcal{L}}_{N}\subset({\mathcal{E}}_{N})^{c},$ (26)
@@
By Lemma 2.13, there exists $c_{1}>0$ such that ${\mathbb{P}}({\mathcal{L}}_{N}^{c})\leq c_{1}^{-1}\exp(-c_{1}n)$, and thus (26) gives
$\mathbb{P}({\mathcal{E}}_{N})\leq\mathbb{P}({\mathcal{E}}^{\prime}_{N})+c_{1}^{-1}\exp(-c_{1}N).$ (27)
To estimate $\mathbb{P}({\mathcal{E}}^{\prime}_{N})$, we first show that the geodesics do not have many closed edges. More precisely, for any $a>0$ there exists $\varepsilon\in(0,1/3)$ such that for all $N\leq\exp(\varepsilon\log^{2}n)$ and $t\geq 1/\varepsilon$,
${\mathbb{P}}(({\mathcal{L}}_{N}^{\prime})^{c})\leq\varepsilon^{-1}\exp(-\varepsilon N),\textrm{ where }{\mathcal{L}}_{N}^{\prime}:=\{\forall\,\gamma\in\cup_{x,y\in\Lambda_{N}}\mathbb{G}(x,y;\Lambda_{tN}):|{\mathrm{clo}}(\gamma)|\leq\tfrac{aN}{\log N}\},$ (28)
where recall that ${\mathrm{clo}}(\gamma)=\{e\in\gamma:e\textrm{ is closed}\}$. Indeed, by Lemmas 2.10 and 2.12 with the remark that $d\geq 3$, there exists $\alpha\in(0,1)$ such that
@@
for all $1/\varepsilon\leq N\leq\exp(\varepsilon\log^{2}n)$ with $\varepsilon=\varepsilon(a,\alpha)$ sufficiently small. Hence, (28) follows.
Suppose now that ${\mathcal{E}}^{\prime}_{N}$ occurs, and let $\pi\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(\Lambda_{tN})$ be a path that realizes ${\mathcal{E}}^{\prime}_{N}$. We assume further that the diameter of $\pi$ is achieved in the first coordinate direction, i.e., there exist $u,v\in\pi$ with $u_{1}-v_{1}=\operatorname{Diam}(\pi)\geq N/2$. Let $\pi^{*}$ be the subpath of $\pi$ from $u$ to $v$. Since $\pi\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(\Lambda_{tN})$, so is $\pi^{*}$. The main idea of the proof can be summarized as follows. We divide the box $\Lambda_{2N}$ into slabs of size $L\asymp\log N$ in the first coordinate. Observe that $\pi^{*}$ has to cross at least $\lfloor N/2L\rfloor$ slabs (since $u_{1}-v_{1}\geq N/2$). By (28), we can assume that $\pi^{*}$ has at most $aN/\log N$ closed edges for some constant $a>0$. Consequently, choosing $L\leq\log N/(5a)$, it follows that $\pi^{*}$ contains at least $\lfloor N/(4L)\rfloor$ open segments inside disjoint slabs, each of diameter at least $L$. For each such segment, the probability that it is not connected to the boundary of $\Lambda_{2N}$ inside its slab is at most $\exp(-cL)$ for some constant $c>0$. Since the slabs are disjoint, these connecting events are independent. Hence, the probability that one of these segments fails to connect to the boundary is at most $\exp(-cN/4)$.
Figure 4. An illustration of the estimate of $\mathbb{P}(\mathcal{E}^{\prime}_{N})$. The black lines represent closed edges, while the red lines represent open edges. The open paths connecting the open segments to $\partial\Lambda_{2N}$ inside their slabs are depicted in blue.
@@
where recall that $u$ is the starting point of $\pi^{*}$. Notice that for all $q\in[K]$, $\pi^{*}$ has to cross the slab $S_{u,q}$ from $H_{u,q-1}$ to $H_{u,q}$. For each $q\in[K]$, let $\pi^{*}_{q}\subset\pi^{*}$ be a subpath of $\pi^{*}\cap S_{u,q}$ such that $\operatorname{Diam}(\pi^{*}_{q})\geq L$. Hence, on the event ${\mathcal{E}}^{\prime}_{N}$,
$\forall q\in[K]:\quad\pi^{*}_{q}\subset S_{u,q},\,\,\operatorname{Diam}(\pi^{*}_{q})\geq L,\,\,\pi^{*}_{q}\not\xleftrightarrow{S_{u,q}}\partial\Lambda_{2N}.$
Let $I:=\{q\in[K]:\pi_{q}^{*}\in{\mathbb{O}}(\Lambda_{N})\}$ be the set of open subpaths of $\pi^{*}$. As $\pi^{*}\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(u,v;\Lambda_{tN})$, there exists $\eta\in\mathbb{G}(\Lambda_{tN})$ such that $\pi^{*}\setminus\eta$ is open. Let $u^{\prime},v^{\prime}$ be the first and last points of $\eta$ intersecting with $\Lambda_{N}$, respectively, ordered from ${\bf s}(\eta)$ to ${\bf e}(\eta)$. Let $\eta^{\prime}$ be the subpath of $\eta$ from $u^{\prime}$ to $v^{\prime}$, so that $\eta^{\prime}\in\cup_{x,y\in\Lambda_{N}}\mathbb{G}(x,y;\Lambda_{tN})$. Since $\pi^{*}\subset\Lambda_{N}$, it follows that $\pi^{*}\setminus\eta=\pi^{*}\setminus\eta^{\prime}$ is open. Then for any $a>0$, conditioned on ${\mathcal{L}}_{N}^{\prime}$ (defined in (28) with $t\geq 1/\varepsilon$ and $\varepsilon=\varepsilon(a)$ sufficiently small), we get that
${\mathrm{clo}}(\pi^{*})\leq{\mathrm{clo}}(\eta^{\prime})\leq\frac{aN}{\log N},$
@@
$|I|\geq K-|{\mathrm{clo}}(\pi^{*})|\geq K-\frac{aN}{\log N}\geq\frac{N}{4L},$ (29)
@@
$\leq\mathbb{P}\Big{(}\exists\,u\in\Lambda_{N},\,\,\exists\,I\subset[K],|I|\geq\tfrac{N}{4L}:\forall q\in I,\,\exists\,\gamma_{q}\in{\mathbb{O}}(\Lambda_{N}),\gamma_{q}\subset S_{u,q},\,\,\operatorname{Diam}(\gamma_{q})\geq L,\gamma_{q}\not\xleftrightarrow{S_{u,q}}\partial\Lambda_{2N}\Big{)}$
$\leq\sum_{u\in\Lambda_{N}}\,\,\sum_{I\subset[K],\,|I|\geq\tfrac{N}{4L}}\mathbb{P}\Big{(}\forall\,q\in I,\exists\,\gamma_{q}\in{\mathbb{O}}(\Lambda_{N}):\operatorname{Diam}(\gamma_{q})\geq L,\gamma_{q}\subset S_{u,q},\gamma_{q}\not\xleftrightarrow{S_{u,q}}\partial\Lambda_{2N}\Big{)}$
$\leq{\mathcal{O}}(1)N^{d}2^{N/L}\Big{(}\mathbb{P}\big{(}\exists\,\gamma\in{\mathbb{O}}(\Lambda_{N}),\gamma\subset S_{0,1}:\operatorname{Diam}(\gamma)\geq L,\gamma\not\xleftrightarrow{S_{0,1}}\partial\Lambda_{2N}\big{)}\Big{)}^{\lfloor\frac{N}{4L}\rfloor},$
where in the last line we have used the independence of events in disjoint slabs. By Lemma 2.14, there exists a constant $C>0$ such that if $L\geq C\log N$,
$\mathbb{P}(\exists\,\gamma\in{\mathbb{O}}(\Lambda_{N})\cap S_{0,1}:\operatorname{Diam}(\gamma)\geq L,\gamma\not\xleftrightarrow{S_{0,1}}\partial\Lambda_{2N})\leq C\exp(-L/C).$
@@
$\mathbb{P}({\mathcal{E}}^{\prime}_{N}\cap{\mathcal{L}}^{\prime}_{N})\leq\exp(-N/(8C)).$ (30)
Finally, we clarify the choice of parameters. We take $a=1/(10C)$ and $L=\lceil C\log N\rceil$ under which $L\leq\log N/(5a)$ and $L\geq C\log N$. Using (30) and (28),
@@
with $\varepsilon=\varepsilon(a)$ given in (28). Combining this with (27) and taking $c_{*}=\frac{1}{4}\min\{\varepsilon,c_{1},1/8C\}$, we obtain the desired result. ∎
@@
Recall the definition of good box in 2.15 with properties (i)–(iii). Fix $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. It follows from Lemmas 2.11 and 2.13 that
$\mathbb{P}(\Lambda_{3N}(e)\text{ does not satisfy (i)})\leq c_{1}^{-1}\exp(-c_{1}N_{\rho}),$ (31)
@@
$\mathbb{P}(\Lambda_{3N}(e)\text{ does not satisfy (ii)})$
$\leq\mathbb{P}(\exists\,x,y\in\Lambda_{3N}(e),\|x-y\|_{\infty}<2N_{\rho},{\mathrm{D}}(x,y)<\infty,{\mathrm{D}}(x,y)\geq 4\rho N_{\rho})\leq c_{2}^{-1}\exp(-c_{2}N).$ (32)
@@
$\mathbb{P}(\Lambda_{3N}(e)\text{ does not satisfy (iii)})$ $\leq\mathbb{P}(\Lambda_{3N}(e)\text{ does not satisfy (iii')})$
$\leq c_{3}^{-1}|{\mathrm{A}_{N}(e)}|^{2}\exp(-c_{3}N_{\rho})\leq c_{3}^{-1}\exp(-c_{3}^{2}N_{\rho}).$ (33)
(iv’) $\text{ there exists }\pi\in{\mathcal{P}}(\Lambda_{3N}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N}(e))\text{ such that }\operatorname{Diam}(\pi)\geq N_{\rho},$
@@
Notice that given a path $\pi\subset\Lambda_{3N}(e)$ with $\operatorname{Diam}(\pi)\geq N_{\rho}$, there exist $x,y\in\pi$ such that $\|x-y\|_{\infty}\geq N_{\rho}$. That is, there exists a subpath $\pi^{\prime}\subset\pi_{x,y}\cap\Lambda_{N_{\rho}/4}(x)$ such that $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$, where $\pi_{x,y}$ is the subpath of $\pi$ from $x$ to $y$. Notice that we always can find $\Lambda\in\mathcal{B}_{N_{\rho}/2}(\Lambda_{3N}(e))$ such that $\Lambda$ contains $\Lambda_{N_{\rho}/4}(x)\cap\Lambda_{3N}(e)$, and hence $\pi^{\prime}\subset\Lambda$ with $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$. Therefore, it follows from (iv’) that there exists a vertex $z\in\Lambda_{3N}(e)$, a crossing cluster ${\mathcal{C}}^{\prime}\subset\Lambda_{N_{\rho}/2}(z)\subset\Lambda_{3N}(e)$, and a subpath $\pi^{\prime}\in{\mathcal{P}}(\Lambda_{N_{\rho}/4}(z))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N_{\rho}/4}(z))$ with $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$, such that $\pi^{\prime}\cap{\mathcal{C}}^{\prime}=\varnothing$. Using Lemma 2.17, there exists $c_{4}>0$ such that for all $N\geq 1$
$\mathbb{P}(\Lambda_{3N}(e)\text{ satisfies (i) but not (iv)})$
@@
Combining this estimate with (31), (2.2.1), and (2.2.1) yields the desired result. ∎
@@
$\{\Lambda_{3N}(e)\text{ is }\textbf{ good}\}\subset\mathcal{V}^{1}_{N}(e)\cap\mathcal{V}^{2}_{N}(e).$ (34)
@@
Next, we assert that $\mathcal{V}^{2}_{N}(e)$ holds. For $x,y\in\Lambda_{3N}(e)$ with ${\mathrm{D}}^{\Lambda_{3N}(e)}(x,y)<\infty$, there exists an open path $\gamma\subset\Lambda_{3N}(e)$ joining $x$ to $y$. Thus, we can find a sequence $(z_{i})_{i=0}^{h}\subset\gamma$ with $h\leq(6N/N_{\rho})^{d}\leq(100\rho^{2})^{d}$ such that $z_{0}=x,z_{h}=y$ and
$\forall i\in[h],\quad\|z_{i-1}-z_{i}\|_{\infty}\leq 2N_{\rho}.$ (35)
Then, by Definition 2.15 (ii), ${\mathrm{D}}^{\Lambda_{4N}(e)}(z_{i-1},z_{i})\leq 4\rho N_{\rho}$. Therefore,
${\mathrm{D}}^{\Lambda_{4N}(e)}(x,y)\leq\sum_{i=1}^{h}{\mathrm{D}}^{\Lambda_{4N}(e)}(z_{i-1},z_{i})\leq h(4\rho N_{\rho})\leq C_{*}N,$
and thus $\mathcal{V}^{2}_{N}(e)$ holds. In summary, we have proved (34) and completed the proof of this lemma. ∎